The emissivity of a spherical body is 0.250. A spherical body with an unknown radius is suspended with a large evacuated enclosure whose walls are at 300K. If the power input required to maintaining the sphere at a temparature of 2500K is 5.5*10^3W, then find the radius of the sphere?   (Hint:-By Stefen-Boltzman Law!!) Get the Answer......

Grade:11

1 Answers

PASUPULETI GURU MAHESH
7 years ago
Here we have   e=0.250 ,Ts=300k,  T=2500  Heat H=5.5×103 and Stefen Boltzman constant =5.67×10-8 w/(m2 k4)
we know that  H= area × e × Stefen Boltzman constant ×(T4-Ts4)
r2=H/(4 pi e × Stefen Boltzman constant ×(T4-Ts4)
you substitute all the values in above and get value .
Thank you